If the current inflation rate is 2.2% and you are earning a real rate of return on investment of 3.8%, then the nominal rate on this investment is closest to (please use the formula; NOT the approximation):        

6.8%        6.08%        6.2%        6.0%
Answer #1

(1+real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.038)=(1+nominal rate)/(1+0.022)

1+nominal rate=(1.038*1.022)

nominal rate=(1.038*1.022)-1

=6.08%(Approx).